# Peng Cheng, and Jiming Chen

### v2

@@ VIII Conclusion and Future Work
In this paper, we introduced Shuffle Differentially Private Policy Elimination (SDP-PE), the first generic algorithm for reinforcement learning under the shuffle privacy model. Our method adapts the policy elimination paradigm to the private learning setting by introducing a novel exponential batching schedule paired with a ”forgetting” mechanism. When instantiated with our proposed Privatizer, SDP-PE achieves a regret of $\widetilde{\mathcal{O}}\left(\sqrt{XAH^{3}K}+\frac{X^{3}AH^{6}}{\varepsilon}\right)$, matching the information-theoretic lower bound up to lower-order terms. Notably, this result achieves utility comparable to the centralized (JDP) model but under a stronger privacy guarantee, while significantly outperforming the local (LDP) model. Our numerical experiments also corroborate these findings and demonstrate the practical advantages of our method. This work thus demonstrates the viability of the shuffle model for achieving a favorable privacy-regret trade-off in data-driven control for advanced CPS. This work opens several avenues for future research. Key directions include improving the computational efficiency of the algorithm for real-time CPS deployment, extending our theoretical guarantees to the setting of function approximation, and exploring the application of alternative privacy notions, such as Rényi DP.
Supporting Lemmas
We first give several supporting lemmas to help analysis.
Lemma 12 (Lemma F.4 in [31])
@@
Let $F_{i}$ for $i=1\cdots$ be a filtration and $X_{1},\cdots,X_{n}$ be a sequence of Bernoulli random variables with $\mathbb{P}\left(X_{i}=1\mid F_{i-1}\right)=P_{i}$ with $P_{i}$ being $F_{i-1}$-measurable and $X_{i}$ being $F_{i}$ measurable. It holds that $\mathbb{P}\left[\exists n:\sum_{t=1}^{n}X_{t}<\sum_{t=1}^{n}P_{t}/2-W\right]\leq e^{-W}.$
Lemma 13 (Lemma 1 in [28])
Let $d>0$ be an integer. Let $\mathcal{X}\subset(\Delta^{d})^{m}$. Then there exists a distribution $\mathcal{D}$ over $\mathcal{X}$, such that $\max_{x=\{x_{i}\}_{i=1}^{dm}\in\mathcal{X}}\sum_{i=1}^{dm}\frac{x_{i}}{y_{i}}=md,$ where $y=\left\{y_{i}\right\}_{i=1}^{dm}=\mathbb{E}_{x\sim\mathcal{D}}[x]$. Moreover, if $\mathcal{X}$ has a boundary set $\partial\mathcal{X}$ with finite cardinality, we can find an approximation solution for $\mathcal{D}$ in poly $(|\partial\mathcal{X}|)$ time.
Proof of Lemma 8
@@
This section provides a detailed proof for bounding the ”Model Bias”—the value difference between the true MDP ($P$) and our absorbing MDP ($P^{\prime}$). The proof proceeds in three main steps. First, we show that our private crude model ($\widetilde{P}^{\text{cru},b}$) is multiplicatively close to the absorbing model for all well-explored tuples. Second, we leverage this to show that for any active policy, the probability of visiting an ”infrequently-visited” tuple is small. Finally, we use this low probability to bound the total difference in value.
@@
Under the absorbing MDP, for any trajectory $S=\{x_{1},a_{1},\dots,x_{h},a_{h}\}$ truncated at time step $h$ such that $(x_{h},a_{a})\in\mathcal{W}$, we have $x_{h^{\prime}}\neq x^{\dagger}$ for any $h^{\prime}\leq h-1$. It holds that
@@
Follow the proof sketch, With these properties established, we can now bound the probability of a trajectory encountering any infrequently-visited tuple (the event $\mathcal{B}$). This is the crucial step in bounding the model bias.
@@
In the first line, $S_{:h+1}$ means the trajectory $S$ truncated at $x_{h+1}$. The second line is because for $(h,x,a,x^{\prime})\notin\mathcal{W}$, $P=P^{\prime}$. The third line is because there is a bijection between trajectories that arrive at $(h,x,a)$ under absorbing MDP and trajectories in $\mathcal{B}_{h}$ that arrive at the same tuple under the original MDP. The last equations follow the definition of $P^{\prime}$ and $\mathcal{B}_{h}$.
@@
To this end, we get the upper bound of the Model Bias term. For any reward function $r^{\prime}$, the non-negative inequality is obvious due to the “absorbing” MDP definition. For the right-hand side, we have that for any policy $\pi\in\phi_{b}$,
@@
[…]
@@
The expectation visitation counts of $(h,x,a)$ from running $\pi^{\text{ref},b}$ for $L_{b}$ episodes is $L_{b}V^{\pi^{\text{ref},b}}(1_{h,x,a},P)\geq L_{b}V^{\pi^{\text{ref},b}}(1_{h,x,a},P^{\prime})$ by definition. According to Lemma 12, with probability $1-\frac{\delta}{HXAK}$, $N_{h}^{\text{ref},b}(x,a)\geq$ $\frac{1}{2}L_{b}V^{\pi^{\text{ref},b}}(1_{h,x,a},P)-\iota\geq\frac{1}{2}L_{b}V^{\pi^{\text{ref},b}}(1_{h,x,a},P^{\prime})-\iota$. The proof is done due to a union bond. ∎
@@
The first inequality is because of the simulation lemma. The second inequality holds by the concentration of $\widetilde{P}^{\text{ref},b}$ by Bernstein inequality and Assumption 4, and Lemma 15. The third inequality is derived via Cauchy-Schwarz inequality, and the property of private counts. The last inequality comes from below.
@@
Proof of Lemma 10
@@
Consider a batch dataset $D$ of $n$ users, we construct private counts $\widetilde{N}_{h}(x,a,x^{\prime}),\widetilde{N}_{h}(x,a),\widetilde{R}_{h}(x,a)$ for all $(h,x,a,x^{\prime})$ by using the proposed Privatizer in in Section V. We prove the two claims of Lemma 10 separately: the privacy guarantee and the utility guarantee.
@@
Privacy guarantee. The proof follows three steps: the adapted shuffle binary summation mechanism is $(\varepsilon^{\prime},\beta^{\prime})$-SDP for one counter, then the Privatizer is $(\varepsilon,\beta)$-SDP for one batch users, and then the entire mechanism for all users is $B$-batch $(\varepsilon,\beta)$-SDP.
@@
Firstly, we consider the shuffle binary mechanism with parameter $(\varepsilon^{\prime},\beta^{\prime})$. Consider two neighboring input $D=(0,d_{1},\dots,d_{n})$ and $D^{\prime}=(1,d_{1},\dots,d_{n})$, we define the random variable $Q$ to be the sum of all the random bits (i.e., $y$ or $y_{1},\dots,y_{m}$) over all users in $D$, and similarly define $Q^{\prime}$ with respect to $D^{\prime}$. We first claim that the output of the shuffler $(\mathcal{F}\circ\mathcal{E}^{n})(D)=Q+\sum_{i=1}^{n}d_{i}$ is $(\varepsilon^{\prime},\beta^{\prime})$-DP. We denote this random mechanism as $M^{*}(\cdot)$.
@@
Define $\Gamma=\sum_{j=1}^{n}d_{j}$ to be the true sum of $D$, and also be the true sum of $D^{\prime}$ minus one. Therefore, for any subset $E\subseteq\mathbb{N}$, we have
@@
Consider we are facing one batch users, for each $(h,x,a)$, we call the shuffle mechanism once. Using the standard composition theorem [5] over all $(h,x,a)$, and fact that one user difference only results in at most $H$ differences in these counters, similar to Lemma 34 of [33], the release of $\{\ddot{N}_{h}(x,a)\}_{h,x,a}$ satisfies $(\frac{\varepsilon}{3},\beta)$-SDP due to the choice of $(\varepsilon^{\prime},\beta^{\prime})$. Similarly, the release of $\{\ddot{N}_{h}(x,a)\}_{h,x,a,x^{\prime}}$ and $\{\widetilde{R}_{h}(x,a)\}_{h,x,a}$ also satisfy $(\frac{\varepsilon}{3},\beta)$-SDP. Then, by post-processing theorem and standard composition theorem in [5] the release of all private counts $\{\widetilde{N}_{h}(x,a)\}_{h,x,a},\{\widetilde{N}_{h}(x,a,x^{\prime})\}_{h,x,a,x^{\prime}},\{\widetilde{R}_{h}(x,a)\}_{h,x,a}$ satisfy $(\varepsilon,\beta)$-SDP. Therefore, by post-processing theorem, for all $B$ batches users, our Privatizer satisfies that $B$-batch $(\varepsilon,\beta)$-SDP.
[…]
Applying the confidence bound on each counter, we can establish the utility bound that with probability $1-3\delta$ for all $(h,x,a,x^{\prime})$,

### v3

@@ IX Conclusion and Future Work
In this paper, we introduce Shuffle Differentially Private Policy Elimination (SDP-PE), the first generic algorithm for reinforcement learning under the shuffle privacy model. Our method successfully adapts the policy elimination paradigm to the private learning setting by introducing a novel exponential batching schedule paired with a “forgetting” mechanism. When instantiated with our proposed Privatizer, SDP-PE achieves a regret of $\widetilde{\mathcal{O}}\left(\sqrt{XAH^{3}K}+\frac{X^{3}AH^{6}}{\varepsilon}\right)$, where the leading term matches the information-theoretic lower bound, and privacy costs are additive. Notably, this result achieves utility comparable to the centralized (JDP) model but under a trustless, distributed and resilient architecture, while significantly outperforming the local (LDP) model. Furthermore, this work unveils a crucial synergy between privacy-driven batching and the reduction of policy deployment costs, decreasing the switching frequency from $\mathcal{O}(K)$ to $\mathcal{O}(\log K)$. This provides a compelling paradigm for data-driven networked systems that demand both privacy and engineering efficiency. Our numerical experiments further corroborate the theoretical findings and demonstrate the practical advantages.
This work opens several promising avenues for future research. Firstly, key directions include improving the computational efficiency of the algorithm and extending its theoretical guarantees to large-scale state-action spaces using function approximation (e.g., deep reinforcement learning). Secondly, it is worth exploring whether it is possible to achieve fewer policy switches (e.g., at the $\Omega(\log\log K)$ level) while satisfying shuffle privacy and maintaining near-optimal regret. Finally, it’s also important to investigate privacy and robustness in more complex networked environments, such as those with malicious clients or more dynamic network topologies.
Appendix A Supporting Lemmas
We first provide several supporting lemmas as follows.
Lemma 12 (Lemma F.4 in [dann2017unifying])
@@ Appendix A Supporting Lemmas
Let $F_{i}$ for $i=1\cdots,n$ be a filtration and $X_{1},\cdots,X_{n}$ be a sequence of Bernoulli random variables with $\mathbb{P}\left(X_{i}=1\mid F_{i-1}\right)=P_{i}$ with $P_{i}$ being $F_{i-1}$-measurable and $X_{i}$ being $F_{i}$ measurable. It holds that $\mathbb{P}\left[\exists k\in[n]:\sum_{t=1}^{k}X_{t}<\sum_{t=1}^{k}P_{t}/2-W\right]\leq e^{-W}.$
Lemma 13 (Lemma 1 in [zhang2022near])
[…]
Appendix B Proof of theoretical results
B-A Proof of Lemma 8
@@
This section provides a detailed proof for bounding the “Model Bias”, the value difference between the true MDP ($P$) and our absorbing MDP ($P^{\prime}$). The proof proceeds in three main steps. First, we show that our private crude model ($\widetilde{P}^{\text{cru},b}$) is multiplicatively close to the absorbing model for all well-explored tuples. Second, we leverage this to show that for any active policy, the probability of visiting an “infrequently-visited” tuple is small. Finally, we use this low probability to bound the total difference in value.
@@
Under the absorbing MDP, for any trajectory $S=\{x_{1},a_{1},\dots,x_{h},a_{h}\}$ truncated at time step $h$ such that $(x_{h},a_{h})\in\mathcal{W}$, we have $x_{h^{\prime}}\neq x^{\dagger}$ for any $h^{\prime}\leq h-1$. It holds that
@@
With these properties established, we now bound the probability of a trajectory encountering any infrequently visited tuple (the event $\mathcal{B}$). This is the crucial step in bounding the model bias.
@@
In the second line, $S_{:h+1}$ means the trajectory $S$ truncated at $x_{h+1}$. The third line is because for $(h,x,a,x^{\prime})\notin\mathcal{W}$, $P=P^{\prime}$. The fourth line is because there is a bijection between trajectories that arrive at $(h,x,a)$ under absorbing MDP and trajectories in $\mathcal{B}_{h}$ that arrive at the same tuple under the original MDP. The last equations follow the definition of $P^{\prime}$ and $\mathcal{B}_{h}$.
@@
Finally, we get the upper bound of the Model Bias term. For any reward function $r^{\prime}$, the non-negative inequality is obvious due to the “absorbing” MDP definition. For the right-hand side, we have that for any policy $\pi\in\phi_{b}$,
@@ B-B Proof of Lemma 9
With probability at least $1-4\delta$, for all $(h,x,a)$, at least one of the following statements holds: (1) For all $x^{\prime}\in\mathcal{X}$, $(h,x,a,x^{\prime})\in\mathcal{W}$, i.e., $P^{\prime}_{h}(x^{\dagger}|x,a)=1$ is fixed and known; (2) $N_{h}^{\text{ref},b}(x,a)\geq\frac{1}{2}L_{b}V^{\pi^{\text{ref},b}}(1_{h,x,a},P^{\prime})-\iota$.
@@
Statement (1) is true by the definition. For statement (2), consider any $(h,x,a)$ tuple, the expectation number of visits to this tuple when running $\pi^{\text{ref},b}$ for $L_{b}$ episodes is $L_{b}V^{\pi^{\text{ref},b}}(1_{h,x,a},P)\geq L_{b}V^{\pi^{\text{ref},b}}(1_{h,x,a},P^{\prime})$ by definition. Thus, according to Lemma 12, with probability $1-\frac{\delta}{HXAK}$, $N_{h}^{\text{ref},b}(x,a)\geq$ $\frac{1}{2}L_{b}V^{\pi^{\text{ref},b}}(1_{h,x,a},P)-\iota\geq\frac{1}{2}L_{b}V^{\pi^{\text{ref},b}}(1_{h,x,a},P^{\prime})-\iota$. A union bound over all such tuples ensures that statement (2) holds with overall high probability. ∎
@@
The first inequality follows the simulation lemma. The second inequality holds by the concentration of $\widetilde{P}^{\text{ref},b}$ by Bernstein inequality, Assumption 4, and Lemma 15. The third inequality is derived via Cauchy-Schwarz inequality, and the property of private counts. The last inequality is obtained by bounding terms $(i)$ and $(ii)$, as detailed below.
@@
B-C Proof of Theorem 6
The proof of Theorem 6 consists of two main parts: bounding the policy switching cost and bounding the cumulative regret. We first give the proof of the upper bound on the policy switching times. The total number of episodes is $K$. The learning process is partitioned into $B$ stages, indexed by $b=1,2,\dots,B$. In each stage $b$, we utilize $3L_{b}$ episodes, where $L_{b}$ episodes are for crude exploration (Algorithm 2) and $2L_{b}$ episodes are for fine exploration (Algorithm 3). The batch lengths are chosen as $L_{b}=2^{b}$, and then the total batch number is $B=\min\{j:3\sum_{b=1}^{j}2^{b}\geq K\}$. Take $j=\lceil\log_{2}(K/6+1)\rceil$, we have $3\sum_{b=1}^{j}2^{b}\geq 3(2^{\log_{2}(K/6+1)}-2)\geq K$. Thus,
$B\leq\log_{2}(K/6+1)+1=\mathcal{O}(\log K).$ (21)
In each stage $b$, policies are switched $H$ times in crude exploration to explore each layer of MDP, and twice in fine exploration. Therefore, the number of policy switching, $N_{switch}$, is bounded by:
$N_{switch}\leq(H+2)B\leq\mathcal{O}(H\log K).$ (22)
We now proceed with bounding the cumulative regret. We first prove the error bound of the value estimate of any policy.
Lemma 19 (Restate of Lemma 7)
With probability $1-9\delta$, it holds that with some constant $C$, for any stage $b$ and $\pi\in\phi_{b}$,
$|V^{\pi}\!(r,\!P)-V^{\pi}\!(\widetilde{r}^{b}\!,\widetilde{P}^{\text{ref},b})|\!\leq\!C\Big(\sqrt{\frac{XAH^{3}\iota}{L_{b}}}+\frac{X^{3}AH^{5}E_{\varepsilon,\delta}\iota}{L_{b}}\Big)\!.$
Proof:
In each stage, we first construct the absorbing MDP $P^{\prime}$, then construct an empirical estimate of $P^{\prime}$, which is $\widetilde{P}^{\text{ref},b}$ as an intermediate term. By applying the triangle inequality, for any $b$ and any $\pi\in\phi^{b}$, we decompose this error into three terms $|V^{\pi}(r,P)-V^{\pi}(\widetilde{r}^{b},\widetilde{P}^{\text{ref},b})|\leq|V^{\pi}(r,P)-V^{\pi}(r,P^{\prime})|+|V^{\pi}(r,P^{\prime})-V^{\pi}(\widetilde{r}^{b},P^{\prime})|+|V^{\pi}(\widetilde{r}^{b},P^{\prime})-V^{\pi}(\widetilde{r}^{b},\widetilde{P}^{\text{ref},b})|$. Combining the bounds from Lemma 8 and Lemma 9, and taking a union bound over all batches $b$, we ensure that with probability at least $1-9\delta$, the value estimation error bound holds for any $b\leq B$ and any $\pi\in\phi^{b}$. ∎
We leverage two key lemmas to prove the regret bound.
Lemma 20
Conditioned on the same high probability event of Lemma 7, the optimal policy $\pi^{*}$ will never be eliminated, i.e., $\pi^{*}\in\phi_{b}$ for $b=1,2,3,\dots,B$.
Proof:
The proof proceeds by induction, since $\phi_{1}$ contains all the deterministic policies, $\pi^{\star}\in\pi_{1}$. Assume $\pi^{\star}\in\phi_{b}$, then we show that
$\sup_{\pi^{\prime}\in\phi_{b}}V^{\pi^{\prime}}(\widetilde{r}^{\text{ref},b},\widetilde{P}^{\text{ref},b})-V^{\pi^{\star}}(\widetilde{r}^{\text{ref},b},\widetilde{P}^{\text{ref},b})$
$=$ $V^{\pi^{b}}(\widetilde{r}^{\text{ref},b},\widetilde{P}^{\text{ref},b})-V^{\pi^{\star}}(\widetilde{r}^{\text{ref},b},\widetilde{P}^{\text{ref},b})$
$\leq$ $\left|V^{\pi^{b}}(\widetilde{r}^{\text{ref},b},\widetilde{P}^{\text{ref},b})\!-\!V^{\pi^{b}}(r,P)\right|\!+\!V^{\pi^{b}}(r,P)\!-\!V^{\pi^{\star}}(r,P)$
$+\left|V^{\pi^{\star}}(r,P)-V^{\pi^{\star}}(\widetilde{r}^{\text{ref},b},\widetilde{P}^{\text{ref},b})\right|$
$\leq$ $2C\left(\sqrt{\frac{XAH^{3}\iota}{L_{b}}}+\frac{X^{3}AH^{5}E_{\varepsilon,\delta}\iota}{L_{b}}\right).$
Thus, the estimated value of the optimal policy $\pi^{\star}$ will not be sufficiently lower than the estimated best policy from $\phi_{b}$. According to the elimination step in Algorithm 1, we will have $\pi^{\star}\in\pi_{b+1}$, which means the optimal policy will never be eliminated. Thus, the optimal policy $\pi^{\star}$ will never be eliminated throughout the entire learning process. ∎
Lemma 21
Conditioned on the same high probability event of Lemma 20, for any remaining policy $\pi\in\phi_{b}$, we have
$\left|V^{\pi^{*}}(r,P)\!-\!V^{\pi}(r,P)\right|\!\leq 4C\left(\sqrt{\frac{XAH^{3}\iota}{L_{b}}}\!+\!\frac{X^{3}AH^{5}E_{\varepsilon,\delta}\iota}{L_{b}}\right).$
Proof:
Since the optimal policy $\pi^{\star}$ is never eliminated, we know that for $\pi\in\phi_{b+1}$, its estimated value must be “close” to that of $\pi^{\star}$. According to the policy elimination rule, it must hold that:
$V^{\pi^{\star}}(\widetilde{r}^{\text{ref},b},\widetilde{P}^{\text{ref},b})-V^{\pi}(\widetilde{r}^{\text{ref},b},\widetilde{P}^{\text{ref},b})$
$\leq$ $\sup_{\pi^{\prime}\in\phi_{b}}V^{\pi^{\prime}}(\widetilde{r}^{\text{ref},b},\widetilde{P}^{\text{ref},b})-V^{\pi}(\widetilde{r}^{\text{ref},b},\widetilde{P}^{\text{ref},b})$
$\leq$ $2C\left(\sqrt{\frac{XAH^{3}\iota}{L_{b}}}+\frac{X^{3}AH^{5}E_{\varepsilon,\delta}\iota}{L_{b}}\right).$
Now, we bound the true sub-optimality gap for any surviving policy $\pi$ by applying the triangle inequality:
$\left|V^{\pi^{*}}(r,P)-V^{\pi}(r,P)\right|$
$\leq$ $\left|V^{\pi^{*}}(r,P)-V^{\pi^{*}}(\widetilde{r}^{\text{ref},b},\widetilde{P}^{\text{ref},b})\right|$
$+\left|V^{\pi^{*}}(\widetilde{r}^{\text{ref},b},\widetilde{P}^{\text{ref},b})-V^{\pi}(\widetilde{r}^{\text{ref},b},\widetilde{P}^{\text{ref},b})\right|$
$+\left|V^{\pi}(\widetilde{r}^{\text{ref},b},\widetilde{P}^{\text{ref},b})-V^{\pi}(r,P)\right|$
$\leq$ $4C\left(\sqrt{\frac{XAH^{3}\iota}{L_{b}}}+\frac{X^{3}AH^{5}E_{\varepsilon,\delta}\iota}{L_{b}}\right).$
Finally, we aggregate the regret over all stages to obtain the total cumulative regret for Theorem 6. The regret for the first stage is at most $3HL_{1}=6H$, because the policies are chosen without much information yet. For stage $b\geq 2$, the policies we use (any policy $\pi\in\phi_{b}$) are at most $4C\left(\sqrt{\frac{XAH^{3}\iota}{L_{b-1}}}+\frac{X^{3}AH^{5}E_{\varepsilon,\delta}\iota}{L_{b-1}}\right)$ sub-optimal, as established by Lemma 21 (note the $L_{b-1}$ as it’s based on the previous stage’s elimination). Therefore, the regret for the $b$-th stage ($3L_{b}$ episodes) is at most $12CL_{b}\left(\sqrt{\frac{XAH^{3}\iota}{L_{b-1}}}+\frac{X^{3}AH^{5}E_{\varepsilon,\delta}\iota}{L_{b-1}}\right)$. The total regret is then bounded by summing over all stages:
$\text{Regret}(K)\!\leq\!6H\!+\!\sum_{b=2}^{B}12CL_{b}\left(\sqrt{\frac{XAH^{3}\iota}{L_{b-1}}}\!+\!\frac{X^{3}AH^{5}E_{\varepsilon,\delta}\iota}{L_{b-1}}\right)$
$\leq 6H+\mathcal{O}\left(\sqrt{XAH^{3}K\iota}\right)+\mathcal{O}\left(X^{3}AH^{5}E_{\varepsilon,\delta}\iota\log K\right)$
$\leq\tilde{\mathcal{O}}\left(\sqrt{XAH^{3}K}+X^{3}AH^{5}E_{\varepsilon,\delta}\right).$
B-D Proof of Lemma 10
@@
Consider a batch dataset $D$ of $n$ users, we construct private counts $\widetilde{N}_{h}(x,a,x^{\prime}),\widetilde{N}_{h}(x,a),\widetilde{R}_{h}(x,a)$ for all $(h,x,a,x^{\prime})$ by using the proposed Privatizer in Section V. We prove the two claims of Lemma 10 separately: the privacy guarantee and the utility guarantee.
@@
Privacy guarantee. The proof follows three steps: the adapted shuffle binary summation mechanism is $(\varepsilon^{\prime},\beta^{\prime})$-SDP for one counter, then the Privatizer is $(\varepsilon,\beta)$-SDP for one batch of users, and then the entire mechanism for all users is $B$-batch $(\varepsilon,\beta)$-SDP.
@@
Firstly, we analyze the shuffle binary mechanism with parameter $(\varepsilon^{\prime},\beta^{\prime})$. Consider two neighboring input $D=(0,d_{1},\dots,d_{n})$ and $D^{\prime}=(1,d_{1},\dots,d_{n})$, we define the random variable $Q$ to be the sum of all the random bits (i.e., $y$ or $y_{1},\dots,y_{m}$) over all users in $D$, and similarly define $Q^{\prime}$ with respect to $D^{\prime}$. We first claim that the output of the shuffler $(\mathcal{F}\circ\mathcal{E}^{n})(D)=Q+\sum_{i=1}^{n}d_{i}$ is $(\varepsilon^{\prime},\beta^{\prime})$-DP. We denote this random mechanism as $M^{*}(\cdot)$.
@@
Let $\Gamma=\sum_{j=1}^{n}d_{j}$ be the true sum of $D$, and also be the true sum of $D^{\prime}$ minus one. Therefore, for any subset $E\subseteq\mathbb{N}$, we have
@@
Next, we address the composition for one batch of users. For each $(h,x,a)$, we invoke the shuffle mechanism once. Using the standard composition theorem [dwork2014algorithmic] over all $(h,x,a)$, and fact that one user difference only results in at most $H$ differences in these counters, similar to Lemma 34 of [hsu2014private], the release of $\{\ddot{N}_{h}(x,a)\}_{h,x,a}$ satisfies $(\frac{\varepsilon}{3},\beta)$-SDP due to the choice of $(\varepsilon^{\prime},\beta^{\prime})$, i.e., $\varepsilon^{\prime}=\frac{\varepsilon}{3H}$ and $\beta^{\prime}=\frac{\beta}{HXA}$. Similarly, the release of $\{\ddot{N}_{h}(x,a)\}_{h,x,a,x^{\prime}}$ and $\{\widetilde{R}_{h}(x,a)\}_{h,x,a}$ also satisfy $(\frac{\varepsilon}{3},\beta)$-SDP. Then, by post-processing theorem and standard composition theorem in [dwork2014algorithmic] the release of all private counts $\{\widetilde{N}_{h}(x,a)\}_{h,x,a},\{\widetilde{N}_{h}(x,a,x^{\prime})\}_{h,x,a,x^{\prime}},\{\widetilde{R}_{h}(x,a)\}_{h,x,a}$ satisfy $(\varepsilon,\beta)$-SDP. Therefore, by post-processing theorem, for all $B$ batches users, our Privatizer satisfies that $B$-batch $(\varepsilon,\beta)$-SDP.
Utility guarantee. Now, we provide the utility guarantee. We first analyze the noise on the true summation. In both regimes of $n$, the output of the mechanism is of the form $Z=\sum_{i=1}^{n}d_{i}+Q-\mathbb{E}[Q]$, and the mechanism is unbiased since $\mathbb{E}[Z-\sum_{i=1}^{n}d_{i}]=\mathbb{E}[Q-\mathbb{E}[Q]]=0$. The error is $Q-\mathbb{E}[Q]$, which is independent of the input $D$ and depends only on the problem parameters. Meanwhile, since in both cases $Q$ is binomial and $\mathbb{E}[Q]\leq\tau$. By Chernoff bounds, we get for any $t>0$, $\mathbb{P}(Q-\mathbb{E}[Q]\leq t)\leq\exp(\frac{-t^{2}}{3\mathbb{E}[Q]})\leq\exp(\frac{-t^{2}}{3\tau})$ and $\mathbb{P}(Q-\mathbb{E}[Q]\geq-t)\leq\exp(\frac{-t^{2}}{3\tau})$. This is the equivalent definition of a sub-Gaussian variable with variance being $\tau=O(\frac{\log(1/\beta)}{{\varepsilon^{\prime}}^{2}})$. Therefore, the error confidence directly follows the concentration of sub-Gaussian variables, i.e., for any $t>0$, we have $\left|Z-\sum_{i=1}^{n}d_{i}\right|>O\left(\frac{\sqrt{\log(1/\beta^{\prime})\log(1/t)}}{\varepsilon^{\prime}}\right)$ with high probability 1t1-t.
Applying the confidence bound to each counter, we can establish the utility bound that with probability $1-3\delta$ for all $(h,x,a,x^{\prime})$,